OpenStudy (anonymous):

find the area of a circle with a radius of 15 millimeters. Round to the nearest tenth.

OpenStudy (anonymous):

Radius = 15 mm = 1.5 cm Area of a circle = Pi*radius*radius = 3.14*(1.5)^2 = 3.14*(2.25) = 7.065 = 7.07